# COMET: Co-Optimization of a CNN Model using Efficient-Hardware OBC Techniques

### v1

@@ I Introduction
Convolutional neural networks (CNNs) have become a cornerstone of modern deep learning, propelled by the rapid progress in artificial intelligence. Their ability to learn features directly from raw data in an end-to-end manner has enabled them to significantly surpass traditional handcrafted methods in key computer vision and pattern recognition tasks, including image classification, object detection, and semantic segmentation [1]. Owing to these advantages, CNNs have found widespread use in diverse domains such as industrial defect inspection, medical imaging, and others [2]. However, despite their success, conventional CNN inference relies on multiplication-intensive operations that are power-hungry and chip-area demanding, posing challenges for edge platforms. To address these challenges, CNNs have been increasingly implemented on FPGAs [3, 4, 5], which are particularly well-suited for CNN acceleration due to their inherent parallelism, reconfigurability, and efficient utilization of on-chip memory and DSP slices [6].
@@
To optimize resource utilization on FPGA, many CNN accelerators adopt the im2col transformation [7] to restructure 2D convolutions into tile-based general matrix multiplication (GEMM) workloads, enabling the reuse of a shared multiply–accumulate (MAC) array across layers [8]. While this strategy improves computational throughput, it exacerbates the demand for DSP multipliers and increases pressure on on-chip memory bandwidth. To mitigate these constraints, various multiplication-reduction techniques have been explored, including Winograd minimal filtering [9], sparse convolution [10], and depthwise separable convolutions [11]. Although these methods effectively reduce the number of operations, they remain tied to DSP-centric architectures—limitations that are particularly pronounced in edge scenarios where both DSP availability and dynamic power budgets are highly constrained.
@@
For efficient CNN implementation on resource-constrained hardware, asymmetric quantization is particularly useful when weight bit-width can be reduced while keeping input bitwidth fixed, albeit at the expense of accuracy [12]. To effectively evaluate such methods, it is essential to select representative models that balance computational complexity and accuracy. Lightweight architectures are particularly well-suited for exploring low-area, low-power design strategies, especially where minimizing energy consumption and silicon footprint is critical. While several works  [13, 14, 4] have targeted complex models such as MobileNetV2 and SqueezeNet—leveraging depthwise separable convolutions and compact network designs to optimize throughput—these approaches often depend on MAC operations implemented through DSP blocks. However, in severely resource-constrained environments, they can become inefficient due to their high computational and energy cost.
@@
LeNet-5 model [15, 16] is often selected as a representative small-scale CNN because of its: (1) simple architecture and low parameter count, which make it suitable for resource-constrained hardware [17]; (2) well-established MNIST performance, providing a reproducible benchmarking baseline [18]; and (3) compatibility with operator types and problem scales typically targeted by GEMM-based frameworks [19]. Unlike the aforementioned models, it avoids architectural complexity, enabling focused studies of arithmetic-level optimizations and low-power, low-area trade-offs. Following this rationale, the present work adopts the classic LeNet-5 CNN as its evaluation benchmark. The model is compact enough to fit entirely on mid-range FPGAs, yet still encompasses representative convolutional, pooling, and fully connected layers. Recent studies [20, 21] consistently demonstrate that accuracy variations introduced by architectural modifications are largely independent of training factors. For instance, replacing the original tanh activations with ReLU leads to better performance in practice [20], while substituting large fully connected layers with global average pooling (GAP) offers a more efficient alternative [21].
@@
Distributed arithmetic (DA) is a multiplier-less approach ideal for low-power, area-constrained hardware [22, 23]. It replaces MAC operations in inner-product computation (IPC) with precomputed look-up tables (LUTs) and a shift-accumulate (SA) unit. This involves representing one vector in two’s complement (TC) [24] or offset-binary coding (OBC) [25], and storing binary or offset binary combinations of the other vector in LUTs. While TC causes exponential LUT growth, OBC reduces this by half through symmetry. Recent works have optimized DA-based LSTMs using hardware-based LUTs [26, 27, 28, 29]. For instance, [26] proposed a parallel OBC-based LUT with pipelining, which reduced critical path delay (CPD) but increased hardware complexity. While [28] improved this with a more efficient approach based on TC. Furthermore, serial LUT designs using high-radix OBC and TC were proposed in [27, 29], incurring additional clock cycle latency. The serial-LUT suffers from limited reusability and significant delay for large dimensions, making pipelining essential but adding latency. To the best of our knowledge, only a few DA-specific implementations targeting either 1-point convolution or CNN acceleration (on ASICs) exist [30, 31, 32]; for example, [30] minimized the LUT size via input pairing, while [31] reduced memory usage through in-memory optimization.
@@
Although various algorithmic optimizations have helped reduce MAC operation counts in CNNs, they still fundamentally rely on DSP-based multipliers, which remain a critical bottleneck for low-power edge deployments. Even advanced dataflow strategies [11, 3] that aim to maximize data reuse fail to eliminate this reliance on MAC units, thereby limiting the efficiency and scalability of accelerators targeting stringent energy and area constraints. This persistent challenge motivates the design of a fully DSP-free CNN architecture based on LeNet-5 model. In this work, we first adopt a streaming (im2col + tiled) strategy to partition the original feature map into multiple small matrix blocks and perform GEMM on each block, thereby avoiding large-scale IPCs in a single pass. To eliminate large pre-computed LUTs, we propose dynamic parallel LUT structure and its variant that enable flexible sharing and reuse via runtime computation of the partial sums. A CNN architecture relying on the OBC scheme is developed and integrated into a general-purpose GEMM core. Using im2col, standard CNN kernels are mapped and executed as OBC-accelerated GEMM operations. We propose COMET (Co-Optimization of a CNN Model using Efficient Hardware OBC Techniques), a framework for efficient CNN acceleration that leverages hardware-optimized OBC techniques to co-optimize performance and resource utilization. The main contributions of this work are as follows:
  • A compact im2col-based formulation of 2D CNN convolution layer using OBC representations of inputs (Scheme A) and weights (Scheme B) enables systematic analysis for hardware optimization.
  • As part of the COMET framework, four novel LUT architectures are introduced using efficient hardware OBC techniques. The shift–accumulate unit is enhanced to merge the pre-scaled bias and offset terms, enabling co-optimization of resource and energy efficiency across multiple clock frequencies.
  • LeNet-5 CNN kernels are mapped onto an OBC-GEMM framework with unified accumulation using Scheme A and Scheme B, and impact of quantization on the accuracy is also analyzed.
  • The design space of the proposed architectures is explored, implemented, and evaluated on an FPGA platform, providing a comprehensive analysis of design trade-offs such as resource utilization and energy efficiency.
@@ II CNN Formulation with OBC
As shown in Fig. 1, a standard CNN is a multi-layer pipeline in which each layer transforms input feature maps of size $H\times W$ into higher-level representations using convolutional kernels of size $K\times L$ and biases. The layer has $C$ input channels and $N$ output channels. Each output channel is obtained by convolving all $C$ input channels with their corresponding kernels via a sliding-window operation (A), summing across channels, and repeating this process for all $N$ outputs. For clarity, we consider the case of batch size $1$, although the formulation extends naturally to multiple batches. The forward pass of a single 2D convolutional layer [33] is
$Y_{n,h,w}=\sum_{c=1}^{C}\sum_{k=1}^{K}\sum_{l=1}^{L}X_{c,h+k,w+l}\ast\Theta_{n,c,k,l}+\beta_{n}$ (1)
where:
  • $X\in\mathbb{R}^{C\times H\times W}$: input tensor (batch size = 1)
  • $\Theta\in\mathbb{R}^{N\times C\times K\times L}$: filter tensor
  • $\beta\in\mathbb{R}^{N}$: bias vector
  • $n$: output channel index $(1\leq n\leq N)$
  • $h,w$: output spatial coordinates
  • $c$: input channel index $(1\leq c\leq C)$
  • $k,l$: kernel indices $(1\leq k\leq K,\;1\leq l\leq L)$
This convolution requires approximately $N\times H\times W\times C\times K\times L$ MAC operations per layer, and across layers, total MACs can reach tens of millions, imposing heavy demands on compute and memory bandwidth [13]. Define
$\tilde{Y}_{n,h,w}=\sum\limits_{c=1}^{C}\sum\limits_{k=1}^{K}\sum\limits_{l=1}^{L}X_{c,h+k,w+l}\ast\Theta_{n,c,k,l}$ (2)
To address this issue, the im2col transformation [7] enables tile-based GEMM, allowing a single MAC array to be reused across layers. Applied to (2), the transformation generally follows these steps:
  • For each output position $(h,w)$, a patch is extracted from the input:
    $\text{patch}_{h,w}(c,k,l)=X_{c,h+k,w+l},\quad\text{with shape }\mathbb{R}^{C\times K\times L}.$
@@
  • Each extracted patch is flattened into a column vector:
    $\text{col}_{X}(h,w)\in\mathbb{R}^{CKL}\triangleq x_{i},\quad\text{with }N\triangleq CKL.$
@@
  • Each filter $\Theta_{n}$ is flattened into a row vector:
    $\text{row}_{\Theta}(n)\in\mathbb{R}^{N}\triangleq\theta_{i}.$
@@
  • The convolution is expressed as an inner product of length $CKL$ between the flattened input patch and filter:
    $\tilde{Y}=\sum_{i=1}^{CKL}\theta_{i}\,x_{i}+\beta_{i}.$ (3)
@@
The resulting output is then reshaped into the appropriate order. For consistency in IPC formulation, we set $C=1$ and $L=1$ for simplicity, while the analysis can be extended for any $C$ and $L$. By denoting the summation in (3) as $\tilde{\tilde{Y}}$, we have
$\tilde{\tilde{Y}}=\sum_{i=1}^{K}\theta_{i}\,x_{i}.$ (4)
@@
In (4), the IPC can be computed by representing either the inputs $x_{i}$ or the weights $\theta_{i}$ in OBC form [23]. For clarity, we refer to these as Scheme A and Scheme B, corresponding to the OBC representation of inputs $x_{i}$ and weights $\theta_{i}$, respectively. For Scheme A, each input $x_{i}$ with bit-width of $B_{1}$-bits is first represented in TC form as
@@
where $b_{i,r}\in\{0,1\}$ is the $r$th bit of $x_{i}$. In OBC representation, $x_{i}$ can be expressed as $x_{i}=\frac{1}{2}[x_{i}-(-x_{i})]$, results (5) in
@@
$\tilde{\tilde{Y}}=\sum_{r=0}^{B_{1}-1}\Bigl(\sum_{i=1}^{K}\tfrac{1}{2}\,\theta_{i}\,\Delta x_{i,r}\Bigr)\,2^{-r}-\Bigl(\tfrac{1}{2}\sum_{i=1}^{K}\theta_{i}\Bigr)\,2^{-(B_{1}-1)}$ (8)
@@
$\tilde{\tilde{Y}}_{r}=\tfrac{1}{2}\sum_{i=1}^{K}\,\theta_{i}\,\Delta x_{i,r}\hskip 5.69046pt\textnormal{and},\hskip 5.69046pt\theta_{ot}=-\tfrac{1}{2}\sum_{i=1}^{K}\theta_{i}$ (9)
@@
$\tilde{\tilde{Y}}=\sum_{r=0}^{B_{1}-1}\tilde{\tilde{Y}}_{r}\,2^{-r}+\theta_{ot}\,2^{-(B_{1}-1)}$ (10)
Substituting (10) into (1) and ignoring the various subscripts, we get
$\tilde{Y}=\sum_{r=0}^{B_{1}-1}\tilde{\tilde{Y}}_{r}\,2^{-r}+(\theta_{ot}+\beta_{i}2^{B_{1}-1})2^{-(B_{1}-1)}$ (11)
@@
$\tilde{Y}=\sum_{r=0}^{B_{1}-1}\tilde{\tilde{Y}}_{r}\,2^{-r}+\tilde{\theta}_{ot}2^{-(B_{1}-1)}$ (12)
@@
$\tilde{Y}=\sum_{r=0}^{B_{2}-1}{\tilde{\tilde{Z}}}_{r}\,2^{-r}+{\tilde{x}}_{ot}2^{-(B_{2}-1)}$ (13)
@@
$\tilde{\tilde{Z}}_{r}$ $=\tfrac{1}{2}\sum_{i=1}^{K}\,x_{i}\,\Delta\theta_{i,r}$
@@ III-A System Overview
Figure 3: System-level diagram of the proposed CNN accelerator.
@@ III-C OBC-GEMM Core
The OBC-GEMM core consists of an array of $L$ LUTs of each size $K$, $L$ SA units, and $L$ PISO units of each size $K$ that handle bit-slices of inputs (or weights), as shown in Fig. 5. The LUTs dynamically generate partial sums using adders and multiplexers for hardware acceleration of IPCs, while SA units efficiently combine these results. PISO units bit-serialize the inputs (or weights) to fit the data flow of the scheme A (or scheme B). A key optimization is a single dedicated adder that computes the offset and bias terms. This eliminates separate bias computation, reducing hardware complexity, and saving area and power.

### v2

@@ I Introduction
Convolutional neural networks (CNNs) have become a cornerstone of modern deep learning. Their ability to learn features directly from raw image data in an end-to-end manner has enabled them to significantly surpass traditional handcrafted methods in key computer vision and pattern recognition tasks, including image classification, object detection, and semantic segmentation [1]. Owing to these advantages, CNNs have found widespread use in diverse domains such as industrial defect inspection, medical imaging, and others [2]. However, despite their success, conventional CNN inference relies on multiplication-intensive operations that are power-hungry and chip-area demanding, posing challenges for edge platforms. To address these challenges, CNNs have been increasingly implemented on FPGAs [3, 4, 5], which are particularly well-suited for CNN acceleration due to their inherent parallelism, reconfigurability, and efficient utilization of on-chip memory and digital signal processing (DSP) slices [6].
@@
To optimize resource utilization on FPGA, many CNN accelerators adopt the im2col transformation [7] to restructure 2D convolutions into tile-based general matrix multiplication (GEMM) workloads, enabling the reuse of a shared multiply–accumulate (MAC) array across layers [8]. While this strategy improves computational throughput, it exacerbates the demand for DSP multipliers and increases pressure on on-chip memory bandwidth. To reduce both the number of DSP multipliers and memory bandwidth, various reduction techniques have been explored, including Winograd minimal filtering [9], sparse convolution [10], and depthwise separable convolutions [11]. Although these methods effectively reduce the number of operations, they remain tied to DSP-centric architectures—limitations that are particularly pronounced in edge scenarios where both DSP availability and dynamic power budgets are highly constrained.
@@
For efficient CNN implementation on resource-constrained hardware, asymmetric quantization is particularly beneficial, as reducing the weight bit-width (while keeping the input bit-width fixed) can significantly lower power and energy consumption, albeit at the expense of some accuracy [12]. To effectively evaluate such methods, it is essential to select representative models that balance computational complexity and accuracy. Lightweight architectures are particularly well-suited for exploring low-area, low-power design strategies, especially where minimizing energy consumption and silicon footprint is critical. While several works  [13, 14, 4] have implemented complex models such as MobileNetV2 and SqueezeNet—leveraging depthwise separable convolutions and compact network designs to optimize throughput—these approaches often depend on MAC operations implemented through DSP blocks. However, in severely resource-constrained environments, they become inefficient due to their high computational and energy cost.
@@
LeNet-5 model [15, 16] is often selected as a representative small-scale CNN because of its: (1) simple architecture and low parameter count, which make it suitable for resource-constrained hardware [17]; (2) well-established MNIST performance, providing a reproducible benchmarking baseline [18]; and (3) compatibility with operator types and problem scales typically targeted by GEMM-based frameworks [19]. Unlike the aforementioned models,Its simple architecture enables a precise study of the implementation trade-offs. Following this rationale, we adopt the classic LeNet-5 CNN as our evaluation benchmark. The model is compact enough to fit entirely on mid-range FPGAs, yet still encompasses representative convolutional, pooling, and fully connected layers. Recent studies [20, 21] indicate that while architectural choices strongly influence model accuracy, their effects are often intertwined with training factors. For example, replacing tanh activations with ReLU has been shown to improve performance in practice [20], and substituting large fully connected layers with global average pooling (GAP) can yield a more efficient design [21].
@@
Distributed arithmetic (DA) is a multiplier-less approach ideal for low-power, area-constrained hardware [22, 23]. It replaces MAC operations in inner-product computation (IPC) with precomputed look-up tables (LUTs) and a shift-accumulate (SA) unit. This involves representing one vector in two’s complement (TC) [24] or offset-binary coding (OBC) [25], and storing binary or offset binary combinations of the other vector in LUTs. While TC causes exponential LUT growth, OBC reduces this by half through symmetry. Recent works have optimized DA-based LSTMs using hardware-based LUTs [26, 27, 28, 29]. For instance, [26] proposed a parallel OBC-based LUT with pipelining, which reduced critical path delay (CPD) but increased hardware complexity. While [28] partially overcomes the limitations of [26] by adopting a TC-based approach, some challenges in optimizing hardware cost remain unaddressed. Furthermore, serial LUT designs using high-radix OBC and TC were proposed in [27, 29], incurring additional clock cycle latency. The serial-LUT suffers from limited reusability and significant delay for large dimensions, making pipelining essential but adding latency. To the best of our knowledge, only a few DA-specific implementations targeting either 1-point convolution or CNN acceleration (on ASICs) exist [30, 31, 32]; for example, [30] minimized the LUT size via input pairing, while [31] reduced memory usage through in-memory optimization.
@@
Although various algorithmic optimizations have helped reduce MAC operation counts in CNNs, they still fundamentally rely on DSP-based multipliers, which remain a critical bottleneck for low-power edge deployments. Even advanced dataflow strategies [11, 3] that aim to maximize data reuse fail to eliminate this reliance on MAC units, thereby limiting the efficiency and scalability of accelerators targeting stringent energy and area constraints. This persistent challenge motivates the design of a fully DSP-free CNN architecture based on LeNet-5 model. In this work, we first adopt a streaming (im2col + tiled) strategy to partition the original feature map into multiple small matrix blocks and perform GEMM on each block, thereby avoiding IPCs on large vectors in a single pass. To eliminate large pre-computed LUTs, we propose four dynamic parallel LUT structures that enable flexible sharing and reuse via runtime computation of the partial sums. A CNN architecture relying on the OBC scheme is developed and integrated into a general-purpose GEMM core. Using im2col, standard CNN kernels are mapped and executed as OBC-accelerated GEMM operations. We name resulting framework COMET (Co-Optimization of a CNN Model using Efficient Hardware OBC Techniques), a framework for efficient CNN acceleration that leverages hardware-optimized OBC techniques to co-optimize performance and resource utilization. Our main contributions are as follows:
  • We propose a compact im2col-based formulation of the 2D CNN convolution layer using OBC representations of inputs (Scheme A) and weights (Scheme B), enabling systematic analysis for hardware optimization.
  • We introduce, as part of the COMET framework, four novel LUT architectures based on efficient hardware OBC techniques. We enhance the shift–accumulate unit to merge pre-scaled bias and offset terms, thereby enabling co-optimization of resource and energy efficiency across multiple clock frequencies.
  • We map LeNet-5 CNN kernels onto an OBC-GEMM framework with unified accumulation using Scheme A and Scheme B, and we analyze the impact of quantization on model accuracy.
  • We explore, implement, and evaluate the design space of the proposed architectures on an FPGA platform, providing a comprehensive analysis of design trade-offs such as resource utilization and energy efficiency.
@@ II CNN Formulation with OBC
As shown in Fig. 1, a standard CNN is a multi-layer pipeline in which each layer transforms input feature maps of size $H\times W$ into higher-level representations using convolutional kernels of size $K\times L$ and biases. A generic layer has $C$ input channels and $N$ output channels. Each output channel is obtained by convolving all $C$ input channels with their corresponding kernels via a sliding-window operation (A), summing across channels, and repeating this process for all $N$ outputs. For clarity, we consider the case of batch size $1$, although the formulation extends naturally to multiple batches. Specifically, given an input tensor with $C$ channels, each with spatial dimensions $H\times W$, i.e., $X\in\mathbb{R}^{C\times H\times W}$, the output of a 2D convolution of $X$ with a filter tensor $\Theta\in\mathbb{R}^{N\times C\times K\times L}$ is a tensor $Y\in\mathbb{R}^{N\times(H+K)\times(W+L)}$, whose entries are given by
$Y_{n,h,w}=\sum_{c=1}^{C}\sum_{k=1}^{K}\sum_{l=1}^{L}X_{c,h+k,w+l}\Theta_{n,c,k,l}+\beta_{n},$ (1)
where $n=1,\ldots,N$, $h=1,\ldots,H+K$, $w=1,\ldots,W+L$, and $\beta_{n}\in\mathbb{R}$ represents a bias term. Note  (1) describes the convolution computation for a single layer. Each product $X_{c,h+k,w+l}\Theta_{n,c,k,l}$ followed by its accumulation constitutes one MAC operation. Therefore, a single convolution layer requires approximately $N\times H\times W\times C\times K\times L$ MAC operations. Across multiple layers, the total MACs can reach tens of millions, imposing heavy demands on compute and memory bandwidth [13]. To address this issue, the im2col transformation [7] enables tile-based GEMM, allowing a single MAC array to be reused across layers. Applied to (1), the transformation generally follows these steps:
  • For each output position $(h,w)$, a 3D patch of the input is extracted as
    $\text{patch}_{h,w}=X[:,\,h:h+K,\,w:w+L]\in\mathbb{R}^{C\times K\times L}.$
@@
  • Each extracted patch $\text{patch}_{h,w}\in\mathbb{R}^{C\times K\times L}$ is flattened into a column vector: Each extracted patch $\text{patch}_{h,w}\in\mathbb{R}^{C\times K\times L}$ is flattened into a column vector
    $\text{col}_{X}(h,w)\in\mathbb{R}^{CKL}\triangleq x_{i},$
    where $N_{p}\triangleq CKL$ denotes the patch dimension.
@@
  • Each filter $\Theta_{n}\in\mathbb{R}^{C\times K\times L}$ is flattened into a row vector:
    $\text{row}_{\Theta}(n)\in\mathbb{R}^{N_{p}}\triangleq\theta_{n}.$
@@
  • The convolution is expressed as an inner product of length $N_{p}=CKL$ between the flattened input patch and filter:
    ${Y}=\sum_{i=1}^{CKL}\theta_{i}\,x_{i}+\beta_{i}.$ (2)
@@
The resulting output is then reshaped into the appropriate order. For consistency in IPC formulation, we set $C=1$ and $L=1$ for simplicity, while the analysis can be extended for any $C$ and $L$. By denoting the summation in (2) as $\tilde{Y}$}, we have
${\tilde{Y}}=\sum_{i=1}^{K}\theta_{i}\,x_{i}.$ (3)
@@
In (3), the IPC in OBC form [23] either the inputs $x_{i}$, a scheme that we refer as Scheme A, or the weights $\theta_{i}$, a scheme that we refer as Scheme B. For Scheme A, each input $x_{i}$ with bit-width of $B_{1}$-bits is first represented in TC form as
@@
where $x_{i,r}\in\{0,1\}$ is the $r$th bit of $x_{i}$. In OBC representation, $x_{i}$ can be expressed as $x_{i}=\frac{1}{2}[x_{i}-(-x_{i})]$, which results (4) in
@@
${\tilde{Y}}=\sum_{r=0}^{B_{1}-1}\Bigl(\sum_{i=1}^{K}\tfrac{1}{2}\,\theta_{i}\,\Delta x_{i,r}\Bigr)\,2^{-r}-\Bigl(\tfrac{1}{2}\sum_{i=1}^{K}\theta_{i}\Bigr)\,2^{-(B_{1}-1)}$ (7)
@@
${\tilde{Y}}_{r}=\tfrac{1}{2}\sum_{i=1}^{K}\,\theta_{i}\,\Delta x_{i,r}\hskip 5.69046pt\textnormal{and},\hskip 5.69046pt\theta_{ot}=-\tfrac{1}{2}\sum_{i=1}^{K}\theta_{i}$ (8)
@@
${\tilde{Y}}=\sum_{r=0}^{B_{1}-1}{\tilde{Y}}_{r}\,2^{-r}+\theta_{ot}\,2^{-(B_{1}-1)}$ (9)
Substituting (9) into (3) and using (2), we get
${Y}=\sum_{r=0}^{B_{1}-1}{\tilde{Y}}_{r}\,2^{-r}+(\theta_{ot}+\beta_{i}2^{B_{1}-1})2^{-(B_{1}-1)}$ (10)
@@
${Y}=\sum_{r=0}^{B_{1}-1}{\tilde{Y}}_{r}\,2^{-r}+\tilde{\theta}_{ot}2^{-(B_{1}-1)}$ (11)
@@
${Y}=\sum_{r=0}^{B_{2}-1}{{\tilde{Z}}}_{r}\,2^{-r}+{\tilde{x}}_{ot}2^{-(B_{2}-1)}$ (12)
@@
${\tilde{Z}}_{r}$ $=\tfrac{1}{2}\sum_{i=1}^{K}\,x_{i}\,\Delta\theta_{i,r}$
@@ III-A System Overview
[…]
@@ III-C OBC-GEMM Core
The OBC-GEMM core consists of an array of $L$ LUTs each of size $K$, $L$ SA units, and $L$ PISO units each of size $K$ that handle bit-slices of inputs (or weights), as shown in Fig. 5. The LUTs dynamically generate partial sums using adders and multiplexers for hardware acceleration of IPCs, while SA units efficiently combine these results. PISO units bit-serialize the inputs (or weights) to fit the data flow of the scheme A (or scheme B). A key optimization is a single dedicated adder that computes the offset and bias terms. This eliminates separate bias computation, reducing hardware complexity, and saving area and power.